Suppose that shoe sizes of American women have a bell-shaped distribution with a mean of 8.48 and a standard deviation of 1.52. Using the empirical rule, what percentage of American women have shoe sizes that are greater than 11.52? Please do not round your answer.
Answer #1

here as z score =(X-mean)/std deviation

therfore percentage of American women have shoe sizes that are greater than 11.52 =P(X>11.52)

=P(Z>(11.52-8.48)/1.52)=P(Z>2) =2.5%